OpenStudy (anonymous):

graph 4^(3x)=8 and i got the answer it is 0,8 is this correct just need to go over answer thanks

OpenStudy (whpalmer4):

\[4^{3x} = 8\]and you think \(x=0.8\)? Did you try substituting that back into the original equation so see if you got a correct number sentence?

OpenStudy (anonymous):

so the answer would be ? i just need to look at the answer to see what i did wrong

OpenStudy (whpalmer4):

\[4^{3x} = 8\]Convert both sides to a power of 2 \[(2*2)^{3x} = 2^3\]\[2^{3x}*2^{3x} = 2^3\]\[2^{6x} = 2^3\]\[6x=3\]\[x=\frac{1}{2}\]

OpenStudy (whpalmer4):

Checking: \[\large 4^{3*\frac{1}{2}} = 4^{\frac{3}{2}} = (\sqrt{4})^3 = 2^3 = 8\checkmark\]

OpenStudy (anonymous):

Thank you!!! it helped alot more than you think.

OpenStudy (anonymous):

so the answer is 8 correct?

OpenStudy (whpalmer4):

no, the answer is \(x=\frac{1}{2}\)

OpenStudy (whpalmer4):

I was verifying that \[4^{3*x} = 8\]when I put \(x = \frac{1}{2}\) in the equation. The result I got was 8, as desired.

OpenStudy (whpalmer4):